What magnitude point charge creates a 10,000 N/C electric field at a distance of .250 m? How large is the field at 10.0 m?
Answer #1

E = k(Q/r^2)

Q=10000*r^2/k

=10000 *(0.250)^2/9*10^9

=6.94*10^8 C


B)
go with proportions


E r^2 = constant;

E1 r1^2 = E2 r2^2

10000 x ( 0.25 )^2 = E2 x (10.0)^2

=> E2 = 10000 x ( 0.25 )^2 / (10.0 )^2 = 6.25 N/C